# Generalizing Medical Image Representations via Quaternion Wavelet Networks

### v3

@@ Abstract
The generalizability challenge in neural network deployment is still a wide-open problem, being amplified by the increasing number of available image datasets coming from different sources and diverse acquisition processes. This matter becomes crucial when dealing with medical image datasets, whose samples are generated by a variety of hospitals, devices, and operators, thus similar datasets can indeed have wide variations. Data of the same medical exam obtained in diverse hospitals may suffer from a domain shifting affecting the generalizability of pretrained models that often have to be re-trained from scratch on the new data [1]. In fact, deep learning techniques have profoundly permeated the field of medical image analysis [2, 3, 4], reaching impressive results in manifold applications, including detection and classification [5, 6, 7], segmentation [8, 9], reconstruction, and inverse problems solution [10, 11, 12], among others. However, such approaches are usually task- or data-specific, thus losing generalizability and undermining their adoption in different contexts. Furthermore, medical images are usually high-quality and full-of-details images that have to be resized and downscaled to be handled by neural networks due to computational limits. This process may compromise image quality and cause a loss of crucial details even in the region of interest, which is often a tiny portion of the whole picture. Due to such image daintiness, the preprocessing often plays a pivotal role in medical analysis [13], being a mandatory step before developing any deep learning model for image datasets [14].
@@
As part of preprocessing, techniques for identifying important and explanatory training examples have been developed for natural images [15, 16] aiming at discarding superfluous data samples and improving the generalization ability of neural networks. Similarly, several works try to establish the most representative patches (or whole sample) to be involved in the training of medical deep learning approaches showing the ability to improve models performance [17, 18]. Simultaneously, a further approach consists of fusing different image modalities in a single and more informative sample [19, 20]. This method is particularly suitable when the dataset contains multimodalities or multiresolution samples. Fusing different modalities combines important information coming from the various acquisitions and may build a more detailed structure, improving pathology examination accuracy. A common approach for medical image fusion aims at leveraging wavelet transforms that split the original image into different frequency components, providing different levels of detail [21, 22]. Recently, a further improvement has been brought by the adoption of the quaternion wavelet transform (QWT) that provides a hypercomplex approach with additional sub-bands and several advantages, including shift invariance, over the standard real-valued wavelet [23, 24, 25, 26].
@@
In this paper, we introduce a novel generalizable framework for medical image analysis and synthesis, referred to as quaternion wavelet network (QUAVE). The proposed approach is data-agnostic, and it does not suffer from domain shifting in data, thus it can be employed with any kind of medical image dataset, for any small and large-scale neural model, and for different tasks. We integrate a quaternion-based preprocessing in a real-valued analysis procedure, to leverage quaternion capabilities in any real-valued scenario, while simultaneously further generalizing the adoption of quaternion and hypercomplex neural networks. QUAVE is built upon two core blocks. The first block exploits quaternion wavelet potentials to extract salient features from medical images that are usually full of details and require detailed preprocessing to preserve crucial content characteristics. These features are then processed by the second QUAVE block, where a neural model extracts the most informative content from the QWT bands to build enhanced representations that can be involved as input to any model for image analysis. Indeed, thanks to the QWT low-frequency and high-frequency sub-bands, we make any model for analysis aware of both approximations and details of the input image, improving its ability to reconstruct or generate images full of details, while also enhancing segmentation capabilities. The overall structure of our framework is depicted in Fig. 1. QUAVE’s output can be generalized to other datasets without requiring the repetition of the training procedure. Furthermore, since the proposed wavelet-based preprocessing is performed in the quaternion domain, we can easily seize quaternion-valued neural networks (QNNs) for the learning stage to further exploit the capabilities of hypercomplex algebra. Indeed, QNNs have been proven to achieve interesting results in processing natural images [27, 28, 29, 30, 31]. Due to the four-dimensional nature of quaternions, QNNs properly handle inputs with dimensions/channels, thus restricting the use of QNNs to a few sets of data. Instead, through QUAVE, we generalize the usage of quaternion and hypercomplex neural networks for single-channel inputs while also improving performance. In fact, by exploiting the proposed approach it is possible to easily handle grayscale and one-channel images in QNNs. Indeed, through the QUAVE processing, we obtain four sub-bands that can be easily encapsulated in a quaternion. Due to the obvious correlations among these sub-bands, QNNs equipped with our method can improve performance by grasping such relations and sharing information among channels.
@@
We prove our theoretical claims through a wide experimental evaluation covering three different tasks and three common medical image datasets. In every test we conduct, QUAVE enhances model performance outperforming previous techniques, according to a plethora of objective metrics and a visual inspection. Therefore, the proposed approach is flexible enough to be adopted with any kind of dataset, especially if it contains images full of details that can be exploited by high-frequency sub-bands.
@@ IV The Proposed QUAVE
In this section, we expound upon our pathway for extracting the most informative representation from the sixteen features obtained through the Quaternion Wavelet Transform (QWT). While the Discrete Wavelet Transform (DWT) produces four sub-bands that can be directly utilized as input for models, the QWT generates sixteen sub-bands, potentially resulting in redundant information if they are used as input to the network. Moreover, quaternion-valued models require exactly four-channel inputs, rendering the direct conveyance of the sixteen QWT sub-bands to the network impractical. To address these challenges, we propose a learning-based approach to build the most informative representation among the sixteen quaternion wavelet sub-bands for involving it as enhanced input to neural models. Recent literature demonstrates the utility of wavelets in various tasks, such as image fusion methods [25] involving multimodal image fusion using wavelets and multiple features. Another area of research utilizes wavelets to enhance super-resolution by enriching output images with improved fine-grain details [46]. Among this class of models, the super-resolution approach comprises a series of image enhancement techniques, including feature refinement [47], which hold significant value for our intended purpose. Based on this premise, we propose a neural model that leverages quaternion wavelets to perform super-resolution, extracting the most informative content by means of the quaternion wavelet features. The underlying idea involves employing transfer learning, wherein QUAVE is trained to perform a super-resolution task on a generic medical dataset (e.g., IXI). Subsequently, the pretrained QUAVE model is utilized to extract the features for a downstream task of interest, such as reconstruction, segmentation, or image translation. Extracting the features entails employing the pretrained QUAVE model and, as depicted in Fig. 4, selecting the weights from the second-last layer, prior to the inverse wavelet transform. The architecture of QUAVE consists of two principal modules. The rough feature extraction module (RFEM) provides the necessary features for performing the quaternion wavelet transform. Subsequently, the wavelet transform module takes the features extracted by the previous modules, it first applies the QWT. Moving forward, the module operates in four groups: one low-frequency group and three high-frequency groups. Before entering the feature refinement block, we reduce the dimensionality of the image features by a factor of four using a convolutional layer. To extract the enriched features of the sub-bands, two linear layers are inserted before the inverse quaternion wavelet transform, which outputs the final super-resolution image. Figure 4 presents an overview of the proposed architecture, which comprises multiple residual groups [48]. Each group consists of residual blocks with convolutional layers and ReLU activation functions. Our methodology allows us to retrieve the features, representing four sub-bands of the quaternion wavelet, that are considered the most informative by the neural model, as they activate the neurons that yield the best performance. During the training phase of a task of interest (e.g. segmentation), we perform inference on pretrained QUAVE, since we utilize the four sub-bands that the model outputs as input for the task of interest. The features representing four sub-bands of the quaternion wavelet can naturally represent a quaternion number, and this fits perfectly for our analysis of this methodology on quaternion-valued neural networks.
@@ V Experimental Setup
We perform experiments in multiple tasks and with different benchmark datasets to show the generalizability of our framework and to strengthen our theoretical claims. To evaluate the effectiveness of our approach, we utilize QUAVE pretrained on a medical grayscale image dataset. We first consider tasks regarding image analysis such as reconstruction and segmentation and then an image synthesis task, that is image modality translation with a large-scale model. We conduct analysis on two algebraic domains: first, we set up experiments to prove that our framework can enhance the performance of real-valued neural networks. Second and most importantly, we perform evaluations with quaternion-valued models to show that QUAVE outperforms other preprocessing techniques and it is flexible to be adopted in different domains, further improving QNNs generalizability. Moreover, we consider diverse types of input images. We take into account medical grayscale and RGB images, and multimodal images, proving that the proposed method can be generalized to any kind of input, thus being data-agnostic and not suffering from domain shifting.
@@ V-A1 Reconstruction
As a first step of our evaluation, we take into account the IXI dataset111https://brain-development.org/ixi-dataset for the reconstruction task. It comprises magnetic resonance images collected from three different hospitals in London, and it is organized as follows: T1, T2, PD-weighted images, MRA images and diffusion-weighted images. We consider the combination of the multimodal T1 and T2 images, namely brain Nuclear Magnetic Resonances (NMR) of and patients respectively. These are provided in NIFTI format, which is a typical format of neuroimaging. In particular, each 3D NRM is composed of different images, each of which identifies a portion of the brain. The sequence of these images is a temporal scan of the area analyzed, so the total number of images in each 3D NRM varies according to the acquisition. For example, in T1 it goes from to , while in T2, it is from to . We preprocess images by keeping only one significant image for the patient into account. After this step, we obtain images from the combination of T1 and T2 ones, each of size . We employ of them for training and for validation.
@@ V-B QUAVE in Combination with Real- and Quaternion-Valued Models
Concerning the segmentation task, we consider a U-Net [51] and a more advanced U-Net++ [52] in our experiments, as they are widespread models for this scenario. The U-Net encoder comprises four blocks, each with two convolutions, batch normalization, and ReLU activation function, while max pooling is interleaved between each block. The number of filters increases going deeper, being , each with kernel sizes and stride . The bottleneck adds a further block with filters. Then, the information is passed to the decoder, which has the same architecture as the encoder except for the replacement of transposed convolutions instead of max pooling operations. A final convolutional layer refines the resulting segmentation map with channel. Similar to the QCAE, the quaternion U-Net (QU-Net) replaces convolutions and transposed convolutions with quaternion-valued counterparts. The more advanced U-Net++ has a similar structure to the original one, inserting however dense connections that aggregate intermediate feature maps. Every hyperparameter is set according to the paper [52]. In these experiments, as for the reconstruction task, we set the learning rate to for the U-Net and to for the U-Net++, as suggested in the paper [52], with Adam optimizer and the batch size to . We train all the models for epochs, repeating the learning process for multiple runs with different seeds.
@@
Regarding the image modality translation task, we consider the recent Target-Aware Generative Adversarial Network (TarGAN) [12]. The TarGAN translates images from a source domain to a target domain while paying particular attention to the region of interest passed as a segmentation map during the translation. The generator takes as input the original image and the four wavelets associated with it, the target area of interest, and the output target modality, generating in the output the translated image in the desired domain. This network comprises two branches, the first one for the whole image and the second one for the region of interest. A single branch is composed of an initial encoder, then a set of blocks shared with the other branch, and finally, a decoder to output the desired image. This network stacks convolutional blocks that employ a convolution layer then the instance normalization, and finally a Leaky ReLU activation function. The upsampling blocks instead involve upsampling layers by interpolation suddenly a convolution operation, and, as in the previous block, finally the instance normalization with Leaky ReLU. The discriminator has a similar structure, for each detail regarding the model’s structure and hyperparameters we refer to the original paper [12]. Overall, this model comprises million parameters, being the largest in our experiments. In this case, we introduce the quaternion-valued counterparts as the quaternion TarGAN has not been introduced yet in literature.

### v5

@@ Abstract
The generalizability challenge in neural network deployment is still a wide-open problem, being amplified by the increasing number of available image datasets coming from different sources and diverse acquisition processes. This matter becomes crucial when dealing with medical image datasets, whose samples are generated by a variety of hospitals, devices, and operators, thus similar datasets can indeed have wide variations. Data of the same medical exam obtained in diverse hospitals may suffer from a domain shifting affecting the generalizability of pretrained models that often have to be re-trained from scratch on the new data [1]. In fact, deep learning techniques have profoundly permeated the field of medical image analysis [2, 3, 4] reaching impressive results in manifold applications, including detection and classification [5, 6, 7], segmentation [8, 9], reconstruction, and inverse problems solution [10, 11, 12], among others. However, such approaches are usually task- or data-specific, thus losing generalizability and undermining their adoption in different contexts. Furthermore, medical images are usually high-quality and full-of-details images that have to be resized and downscaled to be handled by neural networks due to computational limits. This process may compromise image quality and cause a loss of crucial details even in the region of interest, which is often a tiny portion of the whole picture. Due to such image daintiness, the preprocessing often plays a pivotal role in medical analysis [13, 14], being a mandatory step before developing any deep learning model for image datasets [15].
@@
As part of preprocessing, techniques for identifying important and explanatory training examples have been developed for natural images [16, 17] aiming at discarding superfluous data samples and improving the generalization ability of neural networks. Similarly, several works try to establish the most representative patches (or whole sample) to be involved in the training of medical deep learning approaches showing the ability to improve models performance [18, 19]. Simultaneously, a further approach consists of fusing different image modalities in a single and more informative sample [20, 21]. This method is particularly suitable when the dataset contains multimodalities or multiresolution samples. Fusing different modalities combines important information coming from the various acquisitions and may build a more detailed structure, improving pathology examination accuracy. A common approach for medical image fusion aims at leveraging wavelet transforms that split the original image into different frequency components, providing different levels of detail [22, 23, 24]. Recently, a further improvement has been brought by the adoption of the quaternion wavelet transform (QWT) that provides a hypercomplex approach with additional sub-bands and several advantages, including shift invariance, over the standard real-valued wavelet [25, 26, 27, 28].
@@
In this paper, we introduce a novel generalizable framework for medical image analysis and synthesis, referred to as quaternion wavelet network (QUAVE). The proposed approach is data-agnostic, and it does not suffer from domain shifting in data, thus it can be employed with any kind of medical image dataset, for any small and large-scale neural model, and for different tasks. We integrate a quaternion-based preprocessing in a real-valued analysis procedure, to leverage quaternion capabilities in any real-valued scenario, while simultaneously further generalizing the adoption of quaternion and hypercomplex neural networks. QUAVE is built upon two core blocks. The first block exploits quaternion wavelet potentials to extract salient features from medical images that are usually full of details and require detailed preprocessing to preserve crucial content characteristics. These features are then processed by the second QUAVE block, where a neural model extracts the most informative content from the QWT bands to build enhanced representations that can be involved as input to any model for image analysis. Indeed, thanks to the QWT low-frequency and high-frequency sub-bands, we make any model for analysis aware of both approximations and details of the input image, improving its ability to reconstruct or generate images full of details, while also enhancing segmentation capabilities. The overall structure of our framework is depicted in Fig. 1. QUAVE output can be generalized to other datasets without requiring the repetition of the training procedure. Furthermore, since the proposed wavelet-based preprocessing is performed in the quaternion domain, we can easily seize quaternion-valued neural networks (QNNs) for the learning stage to further exploit the capabilities of hypercomplex algebra. Indeed, QNNs have been proven to achieve interesting results in processing images [29, 30, 31, 32, 33, 34, 35]. Due to the four-dimensional nature of quaternions, QNNs properly handle inputs with dimensions/channels, thus restricting the use of QNNs to a few sets of data. Instead, thanks to QUAVE, we generalize the usage of quaternion and hypercomplex neural networks for single-channel inputs while also improving performance. In fact, by exploiting the proposed approach it is possible to easily handle grayscale and one-channel images in QNNs. Indeed, through the QUAVE processing, we obtain four sub-bands that can be easily encapsulated in a quaternion. Due to the obvious correlations among these sub-bands, QNNs equipped with our method can improve performance by grasping such relations and sharing information among channels.
@@
We prove our theoretical claims through a wide experimental evaluation covering three different tasks and three common medical image datasets. In the largest part of the test we conduct, QUAVE enhances model performance outperforming previous techniques, according to a plethora of objective metrics and a visual inspection. Therefore, the proposed approach is flexible enough to be adopted with any kind of dataset, especially if it contains images full of details that can be exploited by high-frequency sub-bands.
@@ IV The Proposed QUAVE
In this section, we expound upon our pathway for extracting the most informative representation from the sixteen features obtained through the Quaternion Wavelet Transform (QWT). While the Discrete Wavelet Transform (DWT) produces four sub-bands that can be directly utilized as input for models, the QWT generates sixteen sub-bands, potentially resulting in redundant information if they are directly used as input to the network. Moreover, quaternion-valued models require exactly four-channel inputs, rendering the direct conveyance of the sixteen QWT sub-bands to the network impractical. To address these challenges, we propose a learning-based approach to build the most informative representation among the sixteen quaternion wavelet sub-bands for involving it as enhanced input to neural models. We build QUAVE upon the recently demonstrated effectiveness of wavelets in various tasks, such as image fusion methods [27] involving multimodal image fusion using wavelets and multiple features. Among these methods, wavelets are particularly effective in enhancing super-resolution by enriching output images with improved fine-grain details [50, 51, 52]. For this reason, we propose a neural model that leverages quaternion wavelets to perform super-resolution, extracting the most informative content by means of the quaternion wavelet features. We propose to leverage transfer learning, wherein QUAVE is trained to perform a super-resolution task on a generic medical dataset (e.g., IXI) and then involved in downstream tasks with the knowledge it has acquired during training.
IV-A Pretraining
We pretrain the model with the objective of enhancing resolution by reconstructing high-quality images with enriched detail through the quaternion wavelet and feature refinement techniques like those in [52]. QUAVE pretraining combines a content loss and a wavelet-based loss component to enforce finer structure and detail preservation in the predictions of the model. The content loss is an L1 loss between the generated image and the ground truth image, which penalizes direct differences in pixel values, while the wavelet loss captures differences in frequency information. This is because the ground truth image and the prediction are transformed into the wavelet domain. The wavelet loss is computed as the L1 loss between these transformed representations and . The final loss function combines the content and wavelet losses, with controlling the influence of the wavelet loss.
Once reached the convergence, the pretrained QUAVE model is utilized to extract the features for a downstream task of interest, such as reconstruction, segmentation, or image translation. Extracting the features entails employing the pretrained QUAVE model and, as depicted in Fig. 4, selecting the weights from the second-last layer, prior to the inverse wavelet transform.
IV-B Architecture
The architecture of QUAVE consists of two principal modules, as shown in Fig. 4. The first one is the rough feature extraction module (RFEM), displayed in Fig. 5, which provides the necessary features for performing the quaternion wavelet transform. The residual groups in RFEM are structured like in [53] and visualized in Fig. 5. Note that to improve the clarity of the figure, the residual connections among the layers are not reported. Subsequently, the wavelet transform module (WTM) takes the features extracted by the RFEM and applies the QWT to those. Moving forward, the module operates in four groups: one low-frequency group and three high-frequency groups. Before entering the feature refinement block, we reduce the dimensionality of the image features by a factor of four using a convolutional layer. Inside the WTM, the Edge Extraction Module focuses on high-frequency information and captures essential edge details of the image. The module utilizes all the high-frequency sub-bands (HL, LH, HH) to create an edge feature map (EFM), which consolidates all high-frequency information to enhance edge detail. The EFM is computed with the L2 Norm:
where represents the high-frequency sub-bands (), and EFM denotes the edge feature map. Next, the edge information is incorporated back into each frequency sub-band through concatenation.
Two linear layers are inserted before the inverse quaternion wavelet transform to extract the enriched features of the sub-bands, which outputs the final super-resolution image. Those layers allow us to retrieve the desired features from QUAVE, representing four sub-bands of the quaternion wavelet, that are considered the most informative by the neural model, as they activate the neurons that yield the best performance.
During the training phase of a downstream task of interest (e.g. segmentation), the original data flows into the pretrained QUAVE that performs inference and outputs the four most representative sub-bands. This enhanced representation is then fed as input to the downstream model to perform the task, as shown in the upper part of Fig. 4. Interestingly, the features representing the four sub-bands of the quaternion wavelet can naturally represent a quaternion number, and this perfectly fits on quaternion-valued neural networks.
@@ V Experimental Setup
We perform experiments in multiple tasks and with different benchmark datasets to show the generalizability of our framework and to strengthen our theoretical claims. Specifically, we utilize QUAVE pretrained on the IXI dataset for downstream tasks involving grayscale medical image datasets and QUAVE pretrained on the Kvasir dataset for downstream tasks on RGB medical image datasets. We first consider tasks regarding image analysis such as reconstruction and segmentation and then an image synthesis task, that is, image modality translation with a large-scale model. We conduct analysis on two algebraic domains: first, we set up experiments to prove that our framework can enhance the performance of real-valued neural networks. When using the term real-valued, we are referring to neural networks defined in the domain of real numbers, therefore not relying on quaternion numbers and thus not involving quaternion-valued layers, to which instead we refer as quaternion-valued. Second and most importantly, we perform evaluations with quaternion-valued models to show that QUAVE outperforms other preprocessing techniques and it is flexible to be adopted in different domains, further improving QNNs generalizability. Moreover, we consider diverse types of input images. We take into account medical grayscale and RGB images, and multimodal images, proving that the proposed method can be generalized to any kind of input, thus being data-agnostic and not suffering from domain shifting.
@@ V-A1 Reconstruction
As a first step of our evaluation, we take into account the IXI dataset111https://brain-development.org/ixi-dataset for the reconstruction task. It comprises magnetic resonance images (MRI) collected from three different hospitals in London, and it is organized as follows: T1, T2, PD-weighted images, MRA images and diffusion-weighted images. We consider the combination of the multimodal T1 and T2 images of and patients respectively. Each 3D MRI scan is stored in NIFTI format and is composed of a stack of 2D slices that together represent the full volumetric structure of the brain. For example, a typical T1-weighted volume comprises approximately 130–150 axial slices, while a T2-weighted volume consists of around 112–136 axial slices. In our experiments, for the purpose of the reconstruction task, we extract a single, representative 2D slice per patient. This approach simplifies the reconstruction task by not processing the entire 3D volume or a large number of slices.
After this step, we obtain images from the combination of T1 and T2 ones, each of size . We employ of them for training and for validation.
@@ V-B QUAVE in Combination with Real- and Quaternion-Valued Models
Concerning the segmentation task, we consider a U-Net [56] and a more advanced U-Net++ [57] in our experiments, as they are widespread models for this scenario. In addition, quaternion counterparts for these architectures are available and have been validated in prior work, which enables us to perform a direct and fair comparison between real-valued and quaternion-valued approaches within our QUAVE framework. Our primary focus is on demonstrating the generalizability and benefits of integrating QUAVE with existing segmentation networks. Thus, using U-Net and U-Net++ allows us to clearly highlight the improvements brought by our framework rather than solely pursuing the highest possible segmentation accuracy. The U-Net encoder comprises four blocks, each with two convolutions, batch normalization, and ReLU activation function, while max pooling is interleaved between each block. The number of filters increases going deeper, being , each with kernel sizes and stride . The bottleneck adds a further block with filters. Then, the information is passed to the decoder, which has the same architecture as the encoder except for the replacement of transposed convolutions instead of max pooling operations. A final convolutional layer refines the resulting segmentation map with channel. Similar to the QCAE, the quaternion U-Net (QU-Net) replaces convolutions and transposed convolutions with quaternion-valued counterparts. The more advanced U-Net++ has a similar structure to the original one, inserting however dense connections that aggregate intermediate feature maps. Every hyperparameter is set according to the paper [57]. In these experiments, as for the reconstruction task, we set the learning rate to for the U-Net and to for the U-Net++, as suggested in the paper [57], with Adam optimizer and the batch size to . We train all the models for epochs, repeating the learning process for multiple runs with different seeds.
| Model | DSC | S-Score | mIoU | MAE |
| TarGAN [12] | ||||
| QUAVE-TarGAN | (+3.4%) | (+10.7%) | (+6.5%) | (+0.4%) |
| QTarGAN | ||||
| QUAVE-QTarGAN | () | (+2.6%) | (+1.6%) | (+0.8%) |
@@
Regarding the image modality translation task, we consider the recent Target-Aware Generative Adversarial Network (TarGAN) [12]. The TarGAN translates images from a source domain to a target domain while paying particular attention to the region of interest passed as a segmentation map during the translation. In the TarGAN framework, segmentation information is explicitly incorporated into the I2I translation process. Specifically, given an image from the source modality and its corresponding target area label , a binarization operation is applied to isolate the target area, yielding , an image containing only the target region. TarGAN then translates both the full input image and the extracted target area into their corresponding representations in the target modality , denoted as . To accomplish this, TarGAN employs a double-stream generator with a shared middle block and two pairs of encoder-decoders: one branch processes the entire image , while the other focuses exclusively on the target area . The generator takes as input the original image and the four wavelets associated with it, the target area of interest, and the output target modality, generating in the output the translated image in the desired domain. This network comprises two branches, the first one for the whole image and the second one for the region of interest. A single branch is composed of an initial encoder, then a set of blocks shared with the other branch, and finally, a decoder to output the desired image. This network stacks convolutional blocks that employ a convolution layer then the instance normalization, and finally a Leaky ReLU activation function. The upsampling blocks instead involve upsampling layers by interpolation suddenly a convolution operation, and, as in the previous block, finally the instance normalization with Leaky ReLU. The discriminator has a similar structure, for each detail regarding the model structure and hyperparameters we refer to the original paper [12]. Overall, this model comprises million parameters, being the largest in our experiments. In this case, we introduce the quaternion-valued counterparts as the quaternion TarGAN has not been introduced yet in literature.
| Model | Params | SSIM | MSE | FID |
| CAE [39] | 305k | |||
| QUAVE-CAE | 305k | (+0%) | (+0%) | (+21.7%) |
| QCAE [39] | 78k | |||
| QUAVE-QCAE | 78k | (+3.5%) | (+33.3%) | (+57.1%) |
| Model | Params | DSC | mIoU | MAE |
| U-Net [56] | 7.8M | |||
| QUAVE-U-Net | 7.8M | (+4.1%) | () | (-2%) |
| U-Net++ [57] | 9.2M | |||
| QUAVE-U-Net++ | 9.2M | (+2.3%) | (+5.4%) | (+7.5%) |
| QU-Net [39] | 1.95M | |||
| QUAVE-QU-Net | 1.95M | (+0.8%) | (+1.4%) | (+3.7%) |
| QU-Net++ [39] | 2.3M | |||
| QUAVE-QU-Net++ | 2.3M | (+0.7%) | (+0.4%) | (+3.1%) |